5.5.31 Expand the expression. If possible, write your answer without exponents. log 3 log3
Answer #1

log, c ) = log (W2-6) - log (W4 ) [iboy g = tega vlogo) = log (22-6)/4 log_(22+4) X5 = balog, (736) – log, (22+4) [:log a=xl